OpenStudy (anonymous):

Can anyone please help with this? How is FOIL used in factoring polynomials?

OpenStudy (anonymous):

(2x-5)(x-4) first: 2x*x= 2x^2 Outside: 2x*(-4)=-8x inside: -5*x=-5x last: (-5)*(-4)=20 2x^2-8x-5x+20 2x^2-13x+20

OpenStudy (anonymous):

Can you explain that in words?

OpenStudy (jim766):

factoring is basically doing foil backwards

OpenStudy (anonymous):

Okay, but how is FOIL used in factoring polynomials?

OpenStudy (jim766):

(x + 2)(x - 3) first foil this x^2 -3x + 2x -6 x^ -1x -6 that is foiling

OpenStudy (anonymous):

Ok..

OpenStudy (jim766):

x^2 - 1x -6 when you factor, you start with a foil solution. ( )( ) now ask yourself , what times what = x^2? (x )(x ) now ask what 2 numbers mult get the last number

OpenStudy (jim766):

(x - 3 ) ( x + 2)

OpenStudy (jim766):

now with factoring....the two numbers need to multiply to get the lastss but add to get the middle number

OpenStudy (jim766):

does that help ?

OpenStudy (anonymous):

Yes, thank you very much.